Differential Analysis Report for Further Processing

The management of Dorsch Aluminum Co. is considering whether to process aluminum ingot further into rolled aluminum. Rolled aluminum can be sold for $4,480 per ton, and ingot can be sold without further processing for $2,508.8 per ton. Ingot is produced in batches of 110 tons by smelting 275 tons of bauxite, which costs $627.2 per ton. Rolled aluminum will require additional processing costs of $940.8 per ton of ingot, and 1.2 tons of ingot will produce 1 ton of rolled aluminum (due to trim losses).

Required:

1. Prepare a report presenting a differential analysis associated with the further processing of aluminum ingot to produce rolled aluminum.

Dorsch Aluminum Co.
Proposal to Process Raw Sugar Further
Differential revenue from further processing per batch:
Revenue from sale of rolled aluminum $_____________
Revenue from sale of aluminum ingot _______________
Differential revenue $______________
Differential cost per batch:
Additional cost of processing rolled aluminum _______________
Differential loss from further processing aluminum ingot, per batch $______________

2. Should Dorsch Aluminum Co. decide to further process aluminum ingot to produce rolled aluminum? Yes or No
Answer #1

Answer 1:

Ingot is produced in batches of 110 tons

1.2 tons of ingot will produce 1 ton of rolled aluminum.

Hence:

Rolled aluminium produced per batch of Ingot = (110 / 1.2) Tons

Revenue from rolled aluminium per batch of ingot = (110 / 1.2) * $4,480 = $410,666.67

Revenue from per batch of Ingot = $2,508.8 * 110 = $275,968

Additional processing costs required to produce rolled aluminum per batch of Ingot = $940.80 * 110 =  $103488

Dorsch Aluminum Co. Proposal to Process Further Differential revenue from further processing per batch: Revenue from sale of

Answer 2:

Yes, Aluminum Co. should decide to further process aluminum ingot to produce rolled aluminum. This will result in earning of additional profit of $31,210.67 per batch of ingot.